OpenStudy (anonymous):

I am trying to figure out my algebra homework with fractions can anyone help me?

OpenStudy (anonymous):

The question is, can you help your self?

OpenStudy (anonymous):

Im 11 years old and really confused and just looking for some help

OpenStudy (anonymous):

come on post the question then

OpenStudy (anonymous):

n/5=7/10 i know the answer is n=7/2 but dont know how they got that answer

OpenStudy (anonymous):

okay, lets work it, you know how to cross multipy?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets do that then: